# MagicLens[Uncaptioned image]: Self-Supervised Image Retrieval with Open-Ended Instructions

### v1

@@ 5.4 Qualitative Study
Figure 8 illustrates top-1 retrieval results on the holdout index set with 1.4M images. Even with complex instruction containing multiple conditions (car and bag examples), MagicLens is still able to accurately comprehend search intents and retrieve desired images. The muffin example showcases that MagicLens can understand the non-trivial temporal relation between images, thanks to the relation diversity introduced by naturally occurring image pairs. However, the image retrieved by MagicLens given the 3D anatomy query may not be generally preferred since the instruction exemplifies the head. This suggests our model may return qualified yet non-perfect examples when the instruction is not clearly expressed. Please refer to Figure 11 in Appendix D for more qualitative studies.
@@ 6 Conclusion
We present MagicLens, a series of self-supervisedly trained image retrieval models that follow open-ended text instructions. Despite being 50$\times$ smaller than prior SOTA methods, MagicLens achieves better results on multiple benchmarks including CIRCO, GeneCIS, and DTIN. Human evaluation on the 1.4M retrieval image pool shows that MagicLens can well satisfy diverse search intents expressed by open-ended instructions, especially complex and beyond visual ones. This indicates MagicLens’ strong capability and potential for real-world search scenarios. Such retrieval models that support open-ended instructions can potentially benefit other vision-language tasks such as visual QA (Antol et al., 2015; Chen et al., 2023c), and enhance multimodal retrieval augmented models (Chen et al., 2022; Hu et al., 2023b). More importantly, we hope our recipe for constructing large-scale synthetic self-supervised training data can shed light on other research directions, such as multimodal retrieval, multimodal representation learning, and beyond.
7 Impact Statements
This work provides novel insights into self-supervised training signals by mining naturally occurring image pairs and develops image retrieval models that follow open-ended instructions to satisfy diverse search intents. It may enable a wide range of search scenarios and have potentials for real-world applications by providing users with more accurate search results. Therefore, we do not see major ethical or social concerns associated with our work.

### v2

@@ 5.4 Qualitative Study
Figure 8 illustrates top-1 retrieval results on the holdout index set with 1.4M images. Even with complex instruction containing multiple conditions (car and bag examples), MagicLens is still able to accurately comprehend search intents and retrieve desired images. The muffin example showcases that MagicLens can understand the non-trivial temporal relation between images, thanks to the relation diversity introduced by naturally occurring image pairs. However, the image retrieved by MagicLens given the 3D anatomy query may not be generally preferred since the instruction exemplifies the head. This suggests our model may return qualified yet imperfect examples when the instruction is not clearly expressed. Please refer to Figure 11 in Appendix D for more qualitative studies.
@@ 6 Conclusion
We present MagicLens, a series of image retrieval models that follow open-ended text instructions. Despite being 50$\times$ smaller than prior SOTA methods, MagicLens achieves better results on multiple benchmarks including CIRCO, GeneCIS, and DTIN. Human evaluation on the 1.4M retrieval image pool shows that MagicLens can well satisfy diverse search intents expressed by open-ended instructions, especially complex and beyond visual ones. This indicates MagicLens’ strong capability and potential for real-world search scenarios. Such retrieval models that support open-ended instructions can potentially benefit other vision-language tasks such as visual QA (Antol et al., 2015; Chen et al., 2023c), and enhance multimodal retrieval augmented models (Chen et al., 2022; Hu et al., 2023b). More importantly, we hope our recipe for constructing large-scale synthetic self-supervised training data can shed light on other research directions, such as multimodal retrieval, multimodal representation learning, and beyond.
Impact Statement
This work provides novel insights into self-supervised training by mining naturally occurring image pairs and develops image retrieval models that follow open-ended instructions to satisfy diverse search intents. It may enable a wide range of search scenarios and have potentials for real-world applications by providing users with more accurate search results. However, despite careful filtering out explicit and offensive images in the training data, MagicLens’ ability to understand image relationships could still be misused for inappropriate image searches. Careful consideration and mitigation strategies are necessary to address these risks.